How many moles of electron weighs one kilogram?
a. 6.023 × 1023
b. 1/9.108× 1031
c. 6.023/9.108× 1054
d. 1/(9.108 ×6.023)× 108

Grade:10

1 Answers

Kevin Nash
9 years ago
Sol. Mass of an electron = 9.108 × 10-31 kg ∵ 9.108 × 10-31 kg = 1.0 electron ∴ 1 kg = 1/〖9.108 ×10〗^(-31) electrons = 〖10〗^31/9.108 × 1/〖6.023 ×10〗^23 = 1/(9.108 ×6.023)× 108 mole of electrons.